# Investigating Non-Transitivity in LLM-as-a-Judge

### v2

@@ 1 Introduction
In this work, we investigate the existence and impact of non-transitivity within AlpacaEval Li et al. (2023), which has been largely overlooked in previous work. AlpacaEval is a popular pairwise comparison framework that employs GPT-4-Turbo as the fixed baseline model. We introduce Soft Non-Transitivity Deviation (SNTD) as a metric to measure the degree of soft non-transitivity in the judge’s continuous preferences and find that LLMs exhibit both hard and soft non-transitive preferences. Additionally, previous studies have demonstrated that LLMs often exhibit various biases Gallegos et al. (2024) such as position bias Zheng et al. (2023); Wang et al. (2023); Zhou et al. (2024), which can lead to spurious correlations in the judge’s preferences. We show that the occurrence of non-transitivity is jointly influenced by position bias and the judge model’s inherent non-transitive reasoning abilities.
@@ 3.1 Measuring Non-Transitivity in Pairwise Comparisons
Estimated Win Rate. We denote the latent quality of the outputs from models $m_{A}$, $m_{B}$, and $m_{C}$ on instruction $I_{i}$ as $\gamma_{A}^{(i)}$, $\gamma_{B}^{(i)}$, and $\gamma_{C}^{(i)}$, respectively. Given empirical observations $\phi$, Bradley-Terry model Bradley & Terry (1952) estimate the quality gap as:
@@ 3.3 Tournament-Based Ranking
Round-Robin Tournament. Tournament-based competition with diverse opponents has been established as an effective approach for performance evaluation in non-transitive games OpenAI et al. (2019); Vinyals et al. (2019), as it enables robust assessment of relative capabilities while mitigating the impact of non-transitivity. Based on this insight, we propose a round-robin tournament structure where each model engages in pairwise evaluation against every other model in the pool, with evaluations conducted by judge $m_{\mathrm{J}}$ over instruction set $\mathcal{I}$. This method enables comprehensive model evaluation through comparisons against a diverse population of models rather than relying on a fixed perspective for assessment. We subsequently apply the BT model to comparison outcomes to assign scores, which are then converted into Elo scores for the final ranking.
Swiss-Wise Iterative Matchmaking Tournament. While round-robin evaluation yields reliable rankings, it presents significant computational challenges at scale. Incorporating a new model into a leaderboard of size $M$ necessitates $(M-1)$ model-level comparisons compared to a single comparison in baseline-fixed frameworks. To address this computational bottleneck, we propose the Swiss-Wise Iterative Matchmaking (Swim) tournament (Algorithm 1), drawing inspiration from binary search and Swiss-system tournaments. Our approach dynamically adjusts matchmaking based on Bradley-Terry coefficients, focusing comparisons near model capability boundaries in a logarithmic manner, thereby reducing the number of comparisons to $\lceil\log_{2}(M)\rceil$.
@@ 4.1 Increased Non-Transitivity with Similar Model
As shown in Table 1, non-transitivity emerges across all four scenarios when GPT-4-Turbo serves as the judge. Both PNT and SNTD generally increase as the performance gap between model pairs $(m_{A},m_{B})$ or $(m_{B},m_{C})$ narrows. Notably, while scenarios LL and ML have identical PNT scores, scenario ML exhibits a higher SNTD value, indicating more non-transitivity. This discrepancy highlights the limitation of the PNT—it fails to capture the continuous nature of judge preferences in assessing non-transitivity.
@@ 4.3 Non-Transitivity is Jointly Influenced by Position Bias and Judge’s Inherent Reasoning Abilities
Prompting Strategies to Mitigate Non-transitivity. We explore various prompting strategies to address non-transitivity in model judgments. Our findings indicate that providing judges with a structured evaluation checklist Cook et al. (2024) marginally reduces non-transitive cases. Notably, while incorporating Chain-of-Thought reasoning Wei et al. (2022) reduces position bias, experimental results indicate that it increases the occurrence of non-transitive preferences. Furthermore, allowing the judge to declare ties not only amplifies position bias but also further increases the likelihood of non-transitivity. See Section C.3 for detailed results.
@@ 6 Limitations and Future Work
Finally, our focus on pairwise comparisons leaves open questions about non-transitivity in pointwise evaluations. While pointwise methods inherently avoid position bias caused by output ordering, converting these scores to pairwise comparison (A > B if score(A) > score(B)) may introduce new forms of non-transitivity, depending on the granularity and consistency of rating criteria. Future work should investigate whether such conversions preserve transitivity and identify conditions that modulate cyclic preferences.
@@ 7 Conclusion
To address the above, we propose a baseline-free framework utilizing round-robin tournaments with Bradley-Terry model, which captures non-transitivity patterns and demonstrates better alignment with human. Recognizing the computational constraints of round-robin tournaments, which require $\mathcal{O}(nm^{2})$ instruction-level comparisons for ranking $m$ models across $n$ instructions, we propose Swim tournaments. This approach achieves $\mathcal{O}(nm\log m)$ complexity through dynamic matching, substantially reducing computational cost while maintaining nearly identical performance.
@@ B.1 Conditions for Non-Transitivity
B.2 Results with Preferences without the Threshold of Ties
@@
Table 3: We measure non-transitivity (without the threshold of ties) in four scenarios, evaluated by GPT-4-Turbo and GPT-3.5-Turbo. Orange cells indicate maximum PNT/SNTD values (highest non-transitivity); blue cells indicate minimum PNT/SNTD values (highest transitivity). When using GPT-4-Turbo as the judge, more non-transitivity can be observed as evaluated model performance becomes more similar and the highest non-transitivity occurs when the performances of all three models are similar; however, GPT-3.5-Turbo does not exhibit this pattern.
@@
To explain the low number of hard non-transitive cases when using GPT-3.5-Turbo as the judge with position switching in Figure 3, we hypothesize that GPT-3.5-Turbo is also affected by other biases, such as verbosity bias Saito et al. (2023) and token bias Alzahrani et al. (2024). Since GPT-3.5-Turbo struggles to accurately assess the quality of outputs, these combined biases influence the judge’s preferences. As a result, even though position switching mitigates the position bias, the averaged preference is still not determined by the actual quality of the outputs but rather by other fixed biases in the prompt, leading to transitive preferences. This observation also motivates us to define the threshold, as it can be used to reduce the impact of other biases.
B.3 Derivation of Expected Win Rate
@@
$\hat{\phi}(o_{A}^{(i)},o_{B}^{(i)}\mid m_{\mathrm{J}},I_{i})=\frac{1}{1+e^{s_{AC}^{(i)}-s_{BC}^{(i)}}}.$ (12)
B.4 Heatmap Implementation
@@
B.5 Preference Distributions of Judge

### v3

@@ 1 Introduction
In this work, we investigate the existence and impact of non-transitivity within AlpacaEval Li et al. (2023), which has been largely overlooked in previous work. AlpacaEval is a popular pairwise comparison framework that employs GPT-4-Turbo as the fixed baseline model. We introduce Soft Non-Transitivity Deviation (SNTD) as a metric to measure the degree of soft non-transitivity in the judge’s continuous preferences and find that LLMs exhibit both hard and soft non-transitive preferences. Additionally, previous studies have demonstrated that LLMs often exhibit various biases Gallegos et al. (2024) such as position bias Zheng et al. (2023); Wang et al. (2024); Zhou et al. (2024b), which can lead to spurious correlations in the judge’s preferences. We show that the occurrence of non-transitivity is jointly influenced by position bias and the judge model’s inherent non-transitive reasoning abilities.
@@ 3.1 Measuring Non-Transitivity in Pairwise Comparisons
Estimated Win Rate. We denote the latent quality of the outputs from models $m_{A}$, $m_{B}$, and $m_{C}$ on instruction $I_{i}$ as $\gamma_{A}^{(i)}$, $\gamma_{B}^{(i)}$, and $\gamma_{C}^{(i)}$, respectively. Given empirical observations $\phi$, Bradley-Terry model estimate the quality gap as:
@@ 3.3 Tournament-Based Ranking
Round-Robin Tournament. Tournament-based competition with diverse opponents has been established as an effective approach for performance evaluation in non-transitive games OpenAI et al. (2019); Vinyals et al. (2019), as it enables robust assessment of relative capabilities while mitigating the impact of non-transitivity. Based on this insight, we propose a round-robin tournament structure where each model engages in pairwise evaluation against every other model in the pool, with evaluations conducted by judge $m_{\mathrm{J}}$ over instruction set $\mathcal{I}$. This method enables comprehensive model evaluation through comparisons against a diverse population of models rather than relying on a fixed perspective for assessment. We subsequently apply the Bradley-Terry model to comparison outcomes to assign scores, which are then converted into Elo scores for the final ranking.
Swiss-Wise Iterative Matchmaking Tournament. While round-robin evaluation yields reliable rankings, it presents significant computational challenges at scale. Incorporating a new model into a leaderboard of size $M$ necessitates $M$ model-level comparisons compared to a single comparison in baseline-fixed frameworks. To address this computational bottleneck, we propose the Swiss-Wise Iterative Matchmaking (Swim) tournament (Algorithm 1), drawing inspiration from binary search and Swiss-system tournaments. Our approach dynamically adjusts matchmaking based on Bradley-Terry coefficients, focusing comparisons near model capability boundaries in a logarithmic manner, thereby reducing the number of comparisons to $\lceil\log_{2}(M)\rceil$.
@@ 4.1 Increased Non-Transitivity with Similar Model
As shown in Table 1, non-transitivity emerges across all four scenarios when GPT-4-Turbo serves as the judge. Both PNT and SNTD generally increase as the performance gap between model pairs $(m_{A},m_{B})$ or $(m_{B},m_{C})$ narrows. Notably, while scenarios LL and ML have identical PNT scores, scenario ML exhibits a higher SNTD value, indicating more non-transitivity. This discrepancy highlights the limitation of the PNT—it fails to capture the continuous nature of judge preferences in assessing non-transitivity. Notably, we observe similar trends across other judges and datasets, confirming the generality of the finding (See Appendix B.2).
@@ 4.3 Non-Transitivity is Jointly Influenced by Position Bias and Judge’s Inherent Reasoning Abilities
Prompting Strategies to Mitigate Non-transitivity. We explore various prompting strategies to address non-transitivity in model judgments. Our analysis focuses on Scenario MM, where the capabilities of the compared models are closely matched, making it easier to observe both non-transitive behaviors and the effects of different prompts. Our findings show that providing judges with a structured evaluation checklist Cook et al. (2024) would marginally reduce non-transitive cases. Interestingly, while incorporating Chain-of-Thought reasoning Wei et al. (2022) helps mitigate position bias, it also leads to a higher incidence of non-transitive preferences. Moreover, allowing the judge to declare ties not only increases position bias but also further amplifies non-transitivity. See Section C.3 for detailed results.
@@ 6 Limitations and Future Work
Secondly, our focus on pairwise comparisons leaves open questions about non-transitivity in pointwise evaluations. While pointwise methods inherently avoid position bias caused by output ordering, converting these scores to pairwise comparison (A > B if score(A) > score(B)) may introduce new forms of non-transitivity, depending on the granularity and consistency of rating criteria. Future work should investigate whether such conversions preserve transitivity and identify conditions that modulate cyclic preferences.
Finally, our analysis relies on the Bradley-Terry model, which assumes transitive model-level preferences by assigning each model a global scalar score. While we do observe instance-level non-transitivity in our pairwise comparisons, these cases are relatively rare, and hard non-transitivity in the aggregated model-level preferences is mild. Therefore, we find the Bradley-Terry model sufficient for our ranking purposes. Nevertheless, we acknowledge that this implementation may not fully capture the nuanced capabilities of models. We leave this as a direction for future work, focusing on more expressive alternatives that parameterize model capabilities in a multi-dimensional space Duan et al. (2017), which remains a promising and under-explored approach for improving the robustness of LLM-as-a-judge evaluations.
@@ 7 Conclusion
To address the above, we propose a baseline-free framework utilizing round-robin tournaments with Bradley-Terry model, which captures non-transitivity patterns and demonstrates better alignment with human. Recognizing the computational constraints of round-robin tournaments, which require $\mathcal{O}(nm^{2})$ instruction-level comparisons for ranking $m$ models across $n$ instructions, we propose Swim tournaments. This approach achieves $\mathcal{O}(nm\log m)$ complexity through dynamic matching, substantially reducing computational cost while maintaining nearly identical performance. The code and data are available at https://github.com/yix8/llm-nontransitivity.
Acknowledgements
We thank to the reviewers and the area chair for their constructive suggestions. We also thank to the OpenAI researcher access program for providing the OpenAI API credits used in this project. Finally, we are grateful to Akbir Khan for early comments, suggestions and advice on the project. LR is supported by the EPSRC Grant EP/S021566/1 and UCL International Scholar Award for Doctoral Training Centres.
@@ B.1 Conditions for Non-Transitivity
B.2 Results Under Varying Judges and Datasets
To further assess the robustness of our findings, we evaluate the same four scenario settings on the AlpacaEval dataset using GPT-4o-mini [^3] Specifically, gpt-4o-mini-2024-07-18 is used for evaluation. as the judge. As shown in Table 3, the results align closely with those obtained using GPT-4-Turbo: the SNTD metric confirms that non-transitivity increases as the performance gap between model pairs narrows. In addition, based on the Chatbot Arena rankings Chiang et al. (2024), GPT-4o-mini is ranked higher than GPT-4-Turbo, suggesting that it serves as a stronger judge. Across almost all scenarios, GPT-4o-mini exhibits lower SNTD and PNT values than GPT-4-Turbo, indicating more transitive judgments. These results provide further empirical support for our claim that stronger judges tend to exhibit less non-transitivity.
To evaluate whether this pattern holds across datasets, we also conduct experiments on the Arena-Hard-Auto Li et al. (2024) dataset, which consists of 500 high-quality prompts curated from Chatbot Arena. Due to computational constraints, we sample 200 prompts for evaluation. We utilize GPT-4-Turbo, GPT-3.5-Turbo, and GPT-4o-mini as judges under the four-scenario framework, selecting models based on their rankings in the Arena-Hard-Auto leaderboard. As shown in the Table 4, the results remain consistent with those observed on AlpacaEval: the SNTD metric confirms that non-transitivity intensifies as the performance gap narrows, particularly for stronger judges. In contrast, GPT-3.5-Turbo exhibits high non-transitivity across all scenarios, due to its inability to reliably distinguish quality differences among the outputs. This consistency suggests that the non-transitive behavior of LLM judges is robust across datasets.
Table 3: We measure non-transitivity on the AlpacaEval dataset across four scenarios, evaluated by GPT-4o-mini. Orange cells indicate maximum PNT/SNTD values (highest non-transitivity); blue cells indicate minimum PNT/SNTD values (highest transitivity). Consistently, more non-transitivity can be observed as evaluated model performance becomes more similar and the highest non-transitivity occurs when the performances of all three models are similar.
Scenarios                       Models GPT-4o-mini
PNT SNTD
[…]
Table 4: We measure non-transitivity on the Arena-Hard-Auto dataset across four scenarios, evaluated by GPT-4-Turbo, GPT-3.5-Turbo, and GPT-4o-mini. Orange cells indicate maximum PNT/SNTD values (highest non-transitivity); blue cells indicate minimum PNT/SNTD values (highest transitivity). We observe a similar pattern as on the AlpacaEval dataset.
Scenarios                       Models GPT-4-Turbo GPT-3.5-Turbo GPT-4o-mini
PNT SNTD PNT SNTD PNT SNTD
[…]
B.3 Results with Preferences without the Threshold of Ties
@@
Table 5: We measure non-transitivity (without the threshold of ties) on the AlpacaEval dataset across four scenarios, evaluated by GPT-4-Turbo and GPT-3.5-Turbo. Orange cells indicate maximum PNT/SNTD values (highest non-transitivity); blue cells indicate minimum PNT/SNTD values (highest transitivity). When using GPT-4-Turbo as the judge, more non-transitivity can be observed as evaluated model performance becomes more similar and the highest non-transitivity occurs when the performances of all three models are similar; however, GPT-3.5-Turbo does not exhibit this pattern.
@@
To explain the low number of hard non-transitive cases when using GPT-3.5-Turbo as the judge with position switching in Figure 5, we hypothesize that GPT-3.5-Turbo is also affected by other biases Zhou et al. (2024a), such as verbosity bias Saito et al. (2023) and token bias Alzahrani et al. (2024). Since GPT-3.5-Turbo struggles to accurately assess the quality of outputs, these combined biases influence the judge’s preferences. As a result, even though position switching mitigates the position bias, the averaged preference is still not determined by the actual quality of the outputs but rather by other fixed biases in the prompt, leading to transitive preferences. This observation also motivates us to define the threshold, as it can be used to reduce the impact of other biases.
B.4 Derivation of Expected Win Rate
@@
$\hat{\phi}(o_{A}^{(i)},o_{B}^{(i)}\mid m_{\mathrm{J}},I_{i})=\frac{1}{1+e^{-\left((\gamma_{A}^{(i)}-\gamma_{C}^{(i)})-(\gamma_{B}^{(i)}-\gamma_{C}^{(i)})\right)}}=\frac{1}{1+e^{-(s_{AC}^{(i)}-s_{BC}^{(i)})}}.$ (11)
B.5 Heatmap Implementation
@@
B.6 Preference Distributions of Judge
@@ Appendix F ELO Scores
We conduct a round-robin tournament to obtain pairwise comparisons and apply the Bradley-Terry model to compute ratings, which are then converted to Elo scores.